OpenStudy (anonymous):

[6.02] Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3

OpenStudy (zepp):

\(4x+y=11\\ y=11-4x\)

OpenStudy (zepp):

substitute y x + 2(11-4x) = 8 x + 22 -8x = 8 -7x +22=8 -7x = 8-22

OpenStudy (precal):

you can always just sub the x and y values to see which ones are true for both equations